OpenStudy (anonymous):

Can anyone just give me a long explanation on how to solve radical expressions with exponent and possibly using an example too?

OpenStudy (anonymous):

I can try ^^

OpenStudy (anonymous):

Ok lets start from something simple.. Say you have something like \[\sqrt{8}\]and we have to simplify... Since the square root of 8 is irritational number it goes on forever with decimals. So putting it in an equation would not be very precise. How do you simplify the square root of 8? Let me explain. Our first step would be to find 2 numbers that multiply together to equal 8 and one of the numbers has to be a perfect square..(i.e 4,9,16,25...) So lets say we pick 4 and 2.. \[\sqrt{4*2}\]

OpenStudy (anonymous):

Ill brb need to go pick someone up real quick itl be like 15 mins

OpenStudy (anonymous):

Okay.

OpenStudy (anonymous):

Ok our equation \[\sqrt{4*2}\] can be simplified even more because 4 is a perfect square. So we split the two apart.. Now we have \[\sqrt{4} \sqrt{2}\] we simplify again and we have \[2\sqrt{2}\] Now this is the simplest form of it. So we are done. Do you understand what i did so far?

OpenStudy (anonymous):

I understand so far. Thanks, and sorry for not responding earlier. I'm very busy right now.